# FSFM: A Generalizable Face Security Foundation Model via Self-Supervised Facial Representation Learning

### v1

@@ Abstract
This work asks: with abundant, unlabeled real faces, how to learn a robust and transferable facial representation that boosts various face security tasks with respect to generalization performance? We make the first attempt and propose a self-supervised pretraining framework to learn fundamental representations of real face images, FSFM, that leverages the synergy between masked image modeling (MIM) and instance discrimination (ID). We explore various facial masking strategies for MIM and present a simple yet powerful CRFR-P masking, which explicitly forces the model to capture meaningful intra-region Consistency and challenging inter-region Coherency. Furthermore, we devise the ID network that naturally couples with MIM to establish underlying local-to-global Correspondence via tailored self-distillation. These three learning objectives, namely 3C, empower encoding both local features and global semantics of real faces. After pretraining, a vanilla ViT serves as a universal vision Foundation Model for downstream Face Security tasks: cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion facial forgery detection. Extensive experiments on 10 public datasets demonstrate that our model transfers better than supervised pretraining, visual and facial self-supervised learning arts, and even outperforms task-specialized SOTA methods.
@@ 1 Introduction
The human face plays a crucial role in daily life and computer vision. In parallel, the facial security landscape suffers from diverse digital and physical manipulations, notably face forgery and presentation attacks. Face forgery edits digital pixels while keeping realistic visual quality. With significant advances in generative models [60, 34, 47], the abuse of this technology, aka deepfakes, could lead to severe trust crises. Presentation attacks use physical media, e.g., printed photos, image/video replays, or 3D masks, to spoof face recognition systems, risking real-life applications like face unlocking and payment. Thus, both academia and industry have been devoted to combating face forgery and presentation attacks with dedicated tasks: Deepfake Detection (DfD), Face Anti-Spoofing (FAS), and recently raised, Diffusion Facial Forgery detection (DiFF).
@@
Current DfD and FAS methods aim to improve generalization to unseen datasets within their respective tasks. DfD methods mainly focus on digital forgery patterns, such as spatiotemporal inconsistency [108, 21, 3], region artifacts [77, 48, 113], and forgery augmentations [113, 77, 3, 89]. FAS methods [50, 134, 62, 133, 92] incorporate domain generalization (DG) techniques to discern physical spoof cues, e.g., paper textures and screen moire patterns. Given these incompatible digital and physical features, most studies treat DfD and FAS as separate face security tasks, i.e., task-specific and hard to transfer. Besides, until now, most DfD and FAS methods still follow fully supervised learning with diverse backbones [22, 95, 28, 43] that are initiated from scratch or supervised ImageNet (natural images) [25] pretraining. However, fully supervised learning requires large-scale annotated data or generative augmentations, which impose expensive resource costs and limit scalability. Moreover, initial weights that lack facial representations may impede capability and generality for face-related tasks [7, 130].
@@
In contrast to supervised learning, self-supervised learning (SSL) takes pretext tasks for pretraining on unlabeled data, where masked image modeling (MIM) [45, 114, 5] and instance discrimination (ID, including contrastive learning [14, 44, 16] and distillation-based [36, 15, 11]) have proven superior performance in various vision tasks. As recent studies [82, 135, 81] suggest that MIM and ID are complementary, SOTA SSL methods [96, 2, 55, 46, 17, 128] combine them to improve the visual representation of natural images. However, SSL progress for facial pretraining, especially security tasks, remains limited, raising Q1: how can face security tasks benefit from self-supervised pretraining to learn scalable and generic representations?
@@
Although existing works introduce SSL to face security tasks, most focus on specific forgery or spoof patterns rather than transferable representations. In DfD, some methods [12, 13, 89, 61] synthesize forgeries from real faces to simulate forgery artifacts like blending, vulnerable to unknown forgery or spoof. Others [31, 41, 127] rely on audio-video datasets and multi-modal information, limiting their scalability. The latest work [123] combines MIM and ID to learn temporal consistency [136] from real face videos, but intra-frame representations fail for image-only forgery and replay videos with consistent temporal features. Conversely, FAS methods employ SSL to mine spoof cues using domain positives [73], domain-invariant semantics [129], or domain alignment [72], yet this FAS domain knowledge contributes less to detecting deepfake and diffusion forgery.
@@
Recent advances in facial pretraining [130, 8, 7, 74, 32] demonstrate that task-agnostic facial representations transfer effectively to several face analysis tasks, such as expression and attribute recognition. However, these methods focus primarily on salient facial features that forgery and spoof faces also exhibit well, i.e., overlook “realness” representations that go beyond normal facial analysis, thus hindering their extrapolation to face security tasks [130]. Furthermore, these methods commonly adopt intra-dataset evaluations for downstream tasks, whereas face security tasks call for cross-dataset generalization. These concerns raise Q2: how to learn universal facial representations from real faces that transfer well to various face security tasks and improve downstream generalization?
@@
To bridge the above gap, instead of task-specific or unified [122, 24, 121] supervised learning, we propose learning the intrinsic properties of unlabeled real face images through SSL. Specifically, we introduce a simple yet effective CRFR-P facial masking strategy into a masked autoencoder [45] that directs attention to pixel-level contexts and region-level correlations, yielding a meaningful and challenging MIM task. By Covering a Random Facial Region (e.g., eyes, nose) and Proportionally masking other regions, our CRFR-P promotes inter-region coherency and intra-region consistency in facial representations. For reliable face-level semantic alignment, we further integrate an ID network with MIM to establish local-to-global correspondence via elaborate self-distillation: the CRFR-P masked online view introduces spatial variance, the unmasked target view retains complete semantics, and Siamese representation decoders form a disentangled space. Thus, our framework, FSFM, leverages the synergy between MIM and ID to empower both local and global perception, learns a universal representation of real faces, transfers well to various face security tasks, and achieves superior generalization in cross-dataset DfD, cross-domain FAS, and unseen DiFF.
Our main contributions are:
@@
We design a simple yet powerful CRFR-P facial masking strategy for MIM to enhance meaningful intra-region Consistency and capture challenging inter-region Coherency, and we develop the ID network coupled with MIM to establish local-to-global Correspondence through elaborate self-distillation—namely, 3C.
@@
We conduct extensive experiments on prevalent face security tasks, including cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion facial forgery detection. Results on 10 public datasets show that our model generalizes better than supervised practices, visual and facial self-supervised pretraining arts, and notably, outperforms task-specialized SOTA methods. Moreover, FSFM benefits from pretraining with more abundantly unlabeled real faces (a free lunch).
@@ 2 Related work
Facial representation learning Several works [93, 97, 79, 100, 66] explore facial SSL to reduce overfitting and improve performance, but they remain tailored for a single task, i.e., task-specific. VLP-based FaRL [130] leverages web image-text CL with MIM to transfer across various facial tasks, but it requires image-text pairs and additional computation overhead, and the text captions tend to describe context rather than face details. Recent SSL efforts aim to learn general facial representations by MIM [8, 103], CL [7, 74, 103], DIS [32, 103] for various downstream tasks, such as expression [8, 7, 74, 32], attribute [8, 32], and AU [7, 74] recognition, alignment [7, 103, 32], etc. Despite excelling in normal face analysis, these methods lack generality for face security tasks as they ignore facial ’realness’ representations.
@@ 3.1 CRFR-P masking
Intuition of intra-region consistency and inter-region coherency. As shown in Fig. 2, random masking and Fasking-I are prone to fully mask small but informative regions (e.g., eyes), hindering accurate learning of textures therein (e.g., pupils, iris). For intra-region consistency, we introduce FRP (Facial Region Proportional) masking that randomly masks an equal portion of patches in each region. FRP ensures all regions have visible patches, thus enabling attention to the same region when restoring masked patches. However, a potential shortcut, restoring a region’s masked patches directly from its unmasked patches, may lead to a trivial reconstruction task and ignore cross-region correlations. In this regard, we formulate CRFR-R (Covering a Random Facial Region followed by Random) masking for inter-region coherency. As the fully masked region can only be derived from visible patches outside it, CRFR-R encourages learning its relationship w.r.t. other regions. Yet, the subsequent random masking may still obscure other small informative regions, undermining the desired consistency within them. FRP and CRFR-R, our preliminary masking strategies, have individual limitations but complement each other.
@@ 3.2 MIM for facial region perception
Online Encoder only operates on the visible patches and maps into latent features , where means the element-wise product for masking and selects the visible ones. The online encoder first embeds by a linear projection as patch embeddings, and adds corresponding positional embeddings . It then processes the fused embeddings through a series of transformer blocks to obtain: .
@@ 3.3 ID for local-to-global self-distillation
Target Encoder takes full patches as the target view to produce target latent features that guide the online encoder in learning holistic representations. Among the target view options in Fig. 3, using full patches enables complete facial semantic embedding for local-to-global representation learning. Thus, the target encoder serves as a teacher that shares the same structure as the student . Similarly, with positional embeddings of full patches, yields global embeddings: .
Online and Target Rep Decoder The online rep decoder maps the full tokens to online representations . Unlike the online decoder , which restores pixel values, reconstructs the masked token representations to align with the uncorrupted target, using significantly shallower transformer blocks followed by a linear layer that predicts features. The token features are output by a mean pooling as the online representations: . In the target branch, the momentum encoder is updated by past iterations of the online encoder , which also serves MIM. If is directly matched with target embeddings , the model may struggle to fit high-level target features while restoring low-level pixels. Thus, a target rep decoder with the same structure as is added to represent the target features in the same disentangled space: .
@@ 4 Experiments
In downstream face security tasks, including deepfake detection (DfD, Sec. 4.2), face anti-spoofing (FAS, Sec. 4.3), and diffusion facial forgery detection (DiFF, Sec. 4.4), we demonstrate the effectiveness of FSFM by asking: Q1: Does our facial representation generalize better than common weight initialization practices? Q2: Is our method superior to other SSL methods for natural vision or normal facial analysis? Q3: Can our pretrained model outperform SOTA specialized methods with only simple fine-tuning? We further ablate FSFM (Sec. 4.5) and present visualizations (Sec. 4.6) to ascertain our contributions. We include more pretraining, finetuning details, ablations, and visualizations in the supplementary material.
@@
| LAA-Net/BI [77] (CVPR’24)‡ | Init(IN) | SD. | 86.28 | 69.69 | 57.13 | 18.3↑ | GS [40] (ICCV’23)‡ | FF++ | 84.97 | 81.65 | 1.8↑ | ||||
| LAA-Net/SBI [77] (CVPR’24)‡ | Init(IN) | SD. | 95.40 | 86.94 | 80.03 | 1.9↑ | UCF [117] (ICCV’23)† | Init(IN) | FF++ | 82.40 | 80.50 | 3.7↑ | |||
@@
| Backbone(or modified): † Xception ‡ EfficientNet-B4 * ViT or transformer-based | |||||||||||||||
@@ 4.1 Pretraining setups and baselines
Pretrained Base Models To answer Q1 and Q2, we include the following pretrained models based on availability (official weights), fairness (vision-only ViT-B/16 backbone), and relevance: 1) ViT-B Scratch [28], to discern pretraining effects versus the backbone itself; 2) ViT-B Sup(IN) [28], supervised ImageNet pretraining, the most common weight initialization for face security tasks; 3) MAE SSL(IN) [45], our MIM network, the baseline across all experiments, including ablations; 4) DINO SSL(IN) [11], a self-distillation method for ID, which also encourages local-to-global correspondence; 5) MCF SSL(LFc) [103], a SOTA facial representation model for face analysis tasks, which also combines MIM and ID and is pretrained on the large-scale LAION-FACE-cropped dataset (20M images). For downstream tasks, we compare our FSFM with these models by initializing the backbone with their pretrained weights and keeping other settings identical. We employ end-to-end fine-tuning for all models, facilitating a fair comparison with SOTA specialized methods to answer Q3.
@@ 4.2 Deepfake detection
| Backbone(or modified): † CNN-based network defined in MADDG ‡ Dense Net # ResNet-18 | ||||||||||||||||
@@
Comparison with SOTA Specialized Methods As shown by Avg.Ours in Tab. 1, FSFM consistently outperforms all counterparts, achieving the best overall performance. Our model surpasses both transformer-based and SSL-based competitors that are task-specific. In particular, our method outperforms SSL-based NACO, a most recent work [123] that also learns consistent representations of real face videos, and it further combines ViT-B with CNN. Our model also performs best across unseen datasets except in video-level testing on CDFV2, where some models trained on synthetic data show slightly better results. These methods (AUNEt, FPG, and LAA-Net with SBIs) simulate pseudo-fake features, especially blending artifacts similar to face-swapping in CDFV2, but struggle with forgeries lacking such clues. Moreover, our method yields notable gains on more challenging DFDC and WDF datasets, which contain diverse unknown real-world manipulations. Remarkably, our FSFM, a vanilla ViT-B base model, achieves SOTA performance without any specialized network modules or tailored data generation for deepfake detection.

### v3

@@ Abstract
This work asks: with abundant, unlabeled real faces, how to learn a robust and transferable facial representation that boosts various face security tasks with respect to generalization performance? We make the first attempt and propose a self-supervised pretraining framework to learn fundamental representations of real face images, FSFM, that leverages the synergy between masked image modeling (MIM) and instance discrimination (ID). We explore various facial masking strategies for MIM and present a simple yet powerful CRFR-P masking, which explicitly forces the model to capture meaningful intra-region Consistency and challenging inter-region Coherency. Furthermore, we devise an ID network that naturally couples with MIM to establish underlying local-to-global Correspondence through tailored self-distillation. These three learning objectives, namely 3C, empower encoding both local features and global semantics of real faces. After pretraining, a vanilla ViT serves as a universal vision Foundation Model for downstream Face Security tasks: cross-dataset deepfake detection, cross-domain face anti-spoofing, and unseen diffusion facial forgery detection. Extensive experiments on 10 public datasets demonstrate that our model transfers better than supervised pretraining, visual and facial self-supervised learning arts, and even outperforms task-specialized SOTA methods.
@@ 1 Introduction
The human face plays a crucial role in daily life and computer vision. In parallel, the facial security landscape suffers from diverse digital and physical manipulations, notably face forgery and presentation attacks. Face forgery edits digital pixels while keeping realistic visual quality. With significant advances in generative models [60, 34, 47], the abuse of such technology, aka deepfakes, has led to severe trust crises. Presentation attacks utilize physical media, e.g., printed photos, image/video replays, or 3D masks, to spoof face recognition systems, risking real-life applications like face unlocking and payment. Thus, both academia and industry have been devoted to combating face forgery and presentation attacks with dedicated tasks: Deepfake Detection (DfD), Face Anti-Spoofing (FAS), and recently raised, Diffusion Facial Forgery detection (DiFF).
@@
Current DfD and FAS methods aim to improve generalization to unseen datasets within their respective tasks. DfD methods mainly focus on digital forgery patterns, such as spatiotemporal inconsistency [109, 21, 3], region artifacts [77, 48, 114], and forgery augmentations [114, 77, 3, 89]. FAS methods [50, 135, 62, 134, 93] incorporate domain generalization techniques to discern physical spoof cues, e.g., paper textures and screen moiré patterns. Given these incompatible digital and physical features, most studies treat DfD and FAS as separate face security tasks, i.e., task-specific and hard-to-transfer. Besides, until now, most DfD and FAS methods still follow fully supervised learning with diverse backbones [22, 96, 28, 43] that are initialized from scratch or supervised ImageNet (natural images) [25] pretraining. However, full supervision requires large-scale annotations or generative augmentations, which incur expensive costs and limit scalability. Moreover, initial weights that lack facial representations may impede capability and generality for face-related tasks [7, 131].
@@
In contrast to supervised learning, self-supervised learning (SSL) takes pretext tasks for pretraining on unlabeled data, where masked image modeling (MIM) [45, 115, 5] and instance discrimination (ID, including contrastive learning [14, 44, 16] and distillation-based [36, 15, 11]) have proven superior performance in various vision tasks. As recent studies [82, 136, 81] suggest that MIM and ID are complementary, SOTA SSL methods [97, 2, 55, 46, 17, 129] combine them to enhance visual representations for natural images. However, SSL progress for facial pretraining, especially security tasks, remains limited, raising Q1: how can face security tasks benefit from self-supervised pretraining to learn scalable and generic representations?
@@
Although existing works introduce SSL to face security tasks, most focus on specific forgery or spoof patterns rather than transferable representations. In DfD, some methods [12, 89, 61, 92] synthesize forgeries from real faces to simulate artifacts like blending, vulnerable to unknown forgery or spoof types. Others [31, 41, 128] rely on audio-video datasets and multimodal information, limiting their scalability. A recent work [124] combines MIM and ID to learn temporal consistency [137] from real face videos, but inter-frame representations fail for image-only forgeries and replay videos with consistent temporal features. Conversely, FAS methods employ SSL to mine spoof cues using domain positives [73], domain-invariant semantics [130], or domain alignment [72], yet this FAS domain knowledge contributes less to detecting deepfake and diffusion forgery.
@@
Recent advances in facial pretraining [131, 8, 7, 74, 32] demonstrate that task-agnostic facial representations effectively transfer to several face analysis tasks, such as expression and attribute recognition. However, these methods focus primarily on salient facial features that forgery and spoof faces also exhibit well, i.e., overlook “realness” representations that go beyond normal facial analysis, thus hindering their extrapolation to face security tasks [131]. Furthermore, these methods typically adopt intra-dataset evaluations for downstream tasks, whereas face security tasks call for cross-dataset generalization. These concerns raise Q2: how to learn universal facial representations from real faces that transfer well to various face security tasks and improve downstream generalization?
@@
To bridge these gaps, instead of task-specific or unified [123, 24, 122] supervised learning, we propose learning the intrinsic properties of unlabeled real face images through SSL. Specifically, we introduce a simple yet effective CRFR-P facial masking strategy into a masked autoencoder [45] that directs attention to pixel-level contexts and region-level correlations, yielding a meaningful and challenging MIM task. By Covering a Random Facial Region (e.g., eyes, nose) and Proportionally masking other regions, our CRFR-P promotes inter-region coherency and intra-region consistency in facial representations. For reliable face-level semantic alignment, we further integrate an ID network with MIM to establish local-to-global correspondence via elaborate self-distillation: the CRFR-P masked online view introduces spatial variance, the unmasked target view preserves complete semantics, and Siamese representation decoders form a disentangled space. Thus, our framework, FSFM, leverages the synergy between MIM and ID to empower both local and global perception, learns a universal representation of real faces, transfers well to various face security tasks, and achieves superior generalization performance. Our main contributions are:
@@
We design a simple yet powerful CRFR-P facial masking strategy for MIM to enhance meaningful intra-region Consistency and capture challenging inter-region Coherency, and we develop the ID network coupled with MIM to establish local-to-global Correspondence through elaborate self-distillation. Namely, 3C.
@@
We conduct extensive experiments on prevalent face security tasks involving 10 datasets: cross-dataset deepfake detection (DfD), cross-domain face anti-spoofing (FAS), and unseen diffusion facial forgery detection (DiFF), which demonstrate our model generalizes better than supervised practices, visual and facial self-supervised pretraining arts, and even outperforms task-specialized SOTA methods. Moreover, FSFM benefits from pretraining with more unlabeled real faces (a free lunch).
@@ 2 Related work
Facial representation learning Several works [94, 98, 79, 101, 66] explore facial SSL to reduce overfitting and improve performance, but they remain tailored for a single task, i.e., task-specific. VLP-based FaRL [131] leverages web image-text CL with MIM to transfer across various facial tasks, but it requires image-text pairs and additional computation overhead, and the text captions tend to describe context rather than face details. Recent SSL efforts aim to learn general facial representations by MIM [8, 104], CL [7, 74, 104], DIS [32, 104] for various downstream tasks, such as expression [8, 7, 74, 32], attribute [8, 32], and AU [7, 74] recognition, alignment [7, 104, 32], etc. Despite excelling in normal face analysis, these methods lack generality for face security tasks as they ignore facial “realness” representations.
@@ 3.1 CRFR-P masking
Intuition of intra-region consistency and inter-region coherency. As shown in Fig. 2, random masking and Fasking-I are prone to fully mask small but informative regions (e.g., eyes), hindering accurate learning of textures therein (e.g., pupils, iris). For intra-region consistency, we introduce FRP (Facial Region Proportional) masking that randomly masks an equal portion of patches in each region. FRP ensures all regions have visible patches, thus enabling attention to the same region when restoring masked patches. However, a potential shortcut, restoring a region’s masked patches directly from its unmasked patches, may lead to a trivial reconstruction task and ignore cross-region correlations. In this regard, we formulate CRFR-R (Covering a Random Facial Region followed by Random) masking for inter-region coherency. As the fully masked region can only be derived from visible patches outside it, CRFR-R encourages learning its relationship w.r.t. other regions. Yet, the subsequent random masking may still obscure other small informative regions, undermining the desired consistency within them. Thus, FRP and CRFR-R, our preliminary masking strategies, have individual limitations but complement each other.
@@ 3.2 MIM for facial region perception
Online Encoder processes only on the visible patches and maps into latent features , where means the element-wise product for masking and selects the visible ones. The online encoder first embeds by a linear projection as patch embeddings, and adds corresponding positional embeddings . It then processes the fused embeddings through a series of transformer blocks to obtain: .
@@ 3.3 ID for local-to-global self-distillation
Target Encoder takes full patches as the target view to produce target latent features that guide the online encoder in learning holistic representations. Among the target view options in Fig. 3, using (c) Full patches enables complete facial semantic embedding for local-to-global representation learning. Thus, the target encoder serves as a teacher that shares the same structure as the student . Similarly, with positional embeddings of full patches, yields global embeddings: .
Online and Target Rep Decoder The online rep decoder maps the full tokens to online representations . Unlike the online decoder , which restores pixel values, reconstructs the masked token representations to align with the uncorrupted target, using significantly shallower transformer blocks followed by a linear layer that predicts features. The token features are output by a mean pooling as the online representations: .
In the target branch, the momentum encoder is updated by past iterations of the online encoder , which also serves MIM. If is directly matched with target embeddings , the model may struggle to fit high-level target features while restoring low-level pixels. Thus, a target rep decoder with the same structure as is added to represent the target features in the same disentangled space: .
@@ 4 Experiments
In downstream face security tasks, including deepfake detection (DfD, Sec. 4.2), face anti-spoofing (FAS, Sec. 4.3), and diffusion facial forgery detection (DiFF, Sec. 4.4), we demonstrate the effectiveness of FSFM by asking: Q1: Does our facial representation generalize better than common weight initialization practices? Q2: Is our method superior to other SSL methods for natural vision or normal facial analysis? Q3: Can our pretrained model outperform SOTA specialized methods with only simple finetuning? We further ablate FSFM (Sec. 4.5) and present visualizations (Sec. 4.6) to ascertain our contributions. We include more pretraining, finetuning details, ablations, and visualizations in the supplementary material.
@@
| LAA-Net/BI [77] (CVPR’24)‡ | Init(IN) | SD | 86.28 | 69.69 | 57.13 | 18.3↑ | GS [40] (ICCV’23)‡ | FF++ | 84.97 | 81.65 | 1.8↑ | ||||
| LAA-Net/SBI [77] (CVPR’24)‡ | Init(IN) | SD | 95.40 | 86.94 | 80.03 | 1.9↑ | UCF [118] (ICCV’23)† | Init(IN) | FF++ | 82.40 | 80.50 | 3.7↑ | |||
@@
| SD-Synthetic (or Self-made) Data Backbone(or modified): † Xception ‡ EfficientNet-B4 * ViT or transformer-based | |||||||||||||||
@@ 4.1 Pretraining setups and baselines
Pretrained Base Models To answer Q1 and Q2, we include the following pretrained models based on availability (official weights), fairness (vision-only ViT-B/16 backbone), and relevance: 1) ViT-B Scratch [28], to discern pretraining effects versus the backbone itself; 2) ViT-B Sup(IN) [28], supervised ImageNet pretraining, the most common weight initialization for face security tasks; 3) MAE SSL(IN) [45], our MIM network, the baseline across all experiments, including ablations; 4) DINO SSL(IN) [11], a self-distillation method for ID, which also encourages local-to-global correspondence; 5) MCF SSL(LFc) [104], a SOTA facial representation model for face analysis tasks, which also combines MIM and ID and is pretrained on the large-scale LAION-FACE-cropped dataset (20M facial images). For downstream tasks, we compare our FSFM with these models by initializing the backbone with their pretrained weights and keeping other settings identical. We employ end-to-end finetuning for all models, facilitating a fair comparison with SOTA specialized methods to answer Q3.
@@ 4.2 Deepfake detection
| DG (Domain Generalization) FAS Technique: DM-Depth Maps AL-Adversarial Learning CL-Contrastive Learning (or triplet, similarity loss) ML-Meta Learning PL-Prototype Learning | ||||||||||||||||
| Backbone(or modified): † CNN-based network defined in MADDG [87] ‡ Dense Net # ResNet-18 | ||||||||||||||||
@@
Comparison with SOTA Specialized Methods As shown by Avg.Ours in Tab. 1, FSFM consistently outperforms all counterparts, achieving the best overall performance. Our model surpasses both transformer-based and SSL-based competitors that are task-specific. In particular, our method outperforms SSL-based NACO, a most recent work [124] that also learns consistent representations of real face videos, and it further combines ViT-B with CNN. Remarkably, our model performs best across unseen datasets except in video-level testing on CDFV2, where some models trained on synthetic data show slightly better results. These methods (AUNet, FPG, and LAA-Net with SBIs) simulate pseudo-fake features, especially blending artifacts similar to face-swapping in CDFV2, but struggle with forgeries lacking such clues. Moreover, our method yields notable gains on more challenging DFDC and WDF datasets, which contain diverse unknown real-world manipulations. Overall, our FSFM, a vanilla ViT-B base model, achieves SOTA performance without any specialized network modules or tailored data generation for deepfake detection.